A hypothesis test for a population proportion p is given below:

Ho: p = 0.25 vs. Ha: p NE 0.25 (NE means not equal)

For sample size n=100 and sample proportion p = 0.30, compute the value of the test statistic:

1.67

-1.12

0.04

1.15
Answer #1

solution:

H0 : p =0.25

Ha : p \neq 0.25

\hat p = 0.30

P0 = 0.25

1 - P0 = 0.75

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.30-0.25/ [\sqrt(0.25*0.75) / 100]

= 1.15